OpenStudy (phagey):

A large bucket that is a full of water has a small leak on the bottom. The bucket loses water at the rate of 0.5 gallons per min. After 6 mins the bucket contains exactly 9 gallons of water. How many minutes does it take for the bucket to lose 7.5 gallons of water? what is the total of minutes it will take for the bucket to be completely empty?

OpenStudy (phagey):

@Michele_Laino can you help me with these last two questions

OpenStudy (michele_laino):

we know that the equation which models such problem is: \(y=12-0.5x\) so, in order to answer to the second part, we have to replqace \(y=0\), so we get: \(0=12-0.5x\) please solve for \(x\)

OpenStudy (phagey):

24

OpenStudy (michele_laino):

correct!

OpenStudy (phagey):

so is that to first question or second

OpenStudy (michele_laino):

it is for second part, after 24 minutes, the bucket will be empty

OpenStudy (phagey):

ok

OpenStudy (michele_laino):

for first part, here is my reasoning: if the bucket lose 0.5 gallons per minutes, then time (in minutes) needed to lose 7.5 gallon, is: \[\Large \frac{{7.5}}{{0.5}} = ...?\]

OpenStudy (michele_laino):

oops.. loses*

OpenStudy (phagey):

is it 15?

OpenStudy (michele_laino):

that's right!

OpenStudy (phagey):

thank you so much u better be glade u real smart in math

OpenStudy (michele_laino):

thanks!! :) :)